OpenStudy (anonymous):

(a + b)(a 2 - ab + b 2)

TheSmartOne (thesmartone):

Noted.

OpenStudy (jdoe0001):

ditto one also may note that \(\textit{difference of cubes} \\ \quad \\ a^3+b^3 = (a+b)(a^2-ab+b^2)\qquad (a+b)(a^2-ab+b^2)= a^3+b^3 \\ \quad \\ a^3-b^3 = (a-b)(a^2+ab+b^2)\qquad (a-b)(a^2+ab+b^2)= a^3-b^3\)